# Equivalence of entropy solutions and gradient flows for pressureless 1D Euler systems

### v2

@@ 1. Introduction
with initial density $\rho(0,x)=\rho_{0}(x)$ and velocity $v(0,x)=v_{0}(x)$, where $f[\rho]$ is a force field generated by the fluid itself. Furthermore, we assume initial data $\rho_{0}\in\mathscr{P}_{2}(\mathbb{R})$, the set of probability measures $\mathscr{P}(\mathbb{R})$ with bounded second moments, and $v_{0}\in L^{2}(\mathbb{R},\rho_{0})$, such that the initial kinetic energy is finite. Moreover, we assume the map $f[\rho]\colon\mathscr{P}_{2}(\mathbb{R})\to\mathscr{M}(\mathbb{R})$, where $\mathscr{M}(\mathbb{R})$ are the signed Borel measures with finite total variation, is absolutely continuous with respect to $\rho$, that is,
@@
We are particularly interested in the two, seemingly distinct, notions of solution introduced in [7] and [36, 37]. Each of which yields a unique solution to their respective systems under appropriate assumptions on the forcing terms, even faced with mass concentration. Our aim is to show that these notions in fact are equivalent. More precisely, we prove the equivalence of Lagrangian solutions [7] and entropy solutions [36, 37] for the Euler system (1.1), which contains the attractive and repulsive Euler–Poisson systems (1.3) as special cases. A very recent paper [43] describes this equivalence between sticky particles flow described in [28] and entropy solutions of (1.1) for $f[\rho]\equiv 0$
@@
with flux function $\mathcal{A}(m)=\alpha(m-m^{2})$ and $M$ being the cumulative distribution function of $\mu$, $M(t,x)=\mu(t,(-\infty,x])$. Here $\alpha=1$ results in an attractive potential, while $\alpha=-1$ gives a repulsive potential. The dynamics of (1.4) is relatively simple in the sense that an attractive potential promotes accumulation of mass $\mu$, while a repulsive potential spreads the mass. This is particularly clear for a sum of Dirac masses $\mu=\sum_{i=1}^{n}m_{i}\delta_{x_{i}}$, i.e., a set of $n$ “particles”, as an attractive potential will aggregate the particles into one particle of unit mass, while a repulsive potential will instantly diffuse the Dirac measures into a measure which is absolutely continuous with respect to the Lebesgue measure. The aggregation and diffusion turn out to have a one-to-one correspondence with respectively shock and rarefaction waves of the conservation law (1.5), see [4, Section 2]. In fact, there the equivalence of (1.4) and (1.5) is first established at the particle level before passing to a limit to cover the general case.
@@
In the above, $\gamma\geq 0$ is a linear damping parameter, while $\tilde{\phi}=\frac{1}{2}x^{2}-\lvert x\rvert$ represents a potential with Coulomb repulsion and quadratic confinement. This system has been studied in detail for in [14] for classical solutions up until blow-up, or mass concentration, for which precise critical thresholds are given. We want to use the aforementioned solution concepts to study (1.6) beyond the time of blow-up and its asymptotic behavior.
@@ 1.1. Main results
(1.8) $\frac{v(t,x_{2})-v(t,x_{1})}{x_{2}-x_{1}}\leq\frac{1}{t}\enspace\text{for $\rho$-a.e.}\ x_{1}\leq x_{2}.$
@@
[…]
@@ 2.1. First considerations and simplifications
The first term on the right-hand side of (2.2), scaled by $\alpha$, is the force per unit mass coming from the Poisson interaction forces between particles, and is therefore an internal force for the system. On the other hand, the second term scaled by $\beta$ is the force per unit mass coming from an external force, which could be an external gravitational or electrical field. Introducing the kinetic energy $E_{\mathrm{k}}(\{x_{i}\})=\frac{1}{2}\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{2}$, then the Euler–Lagrange equations for the Lagrangian $L=E_{\mathrm{k}}-E_{\mathrm{p}}$ yields (2.2). Moreover, the energy $E_{\mathrm{k}}+E_{\mathrm{p}}$, which corresponds to the Hamiltonian after introducing the momentum $p_{i}=m_{i}\dot{x}_{i}$, is conserved. The total momentum $\sum_{i=1}^{n}m_{i}\dot{x}_{i}$ however, is not conserved because of the external force. If we introduce the center of mass $\bar{x}\coloneqq\sum_{i=1}^{n}m_{i}x_{i}$ and average velocity, which here coincides with the total momentum, $\bar{v}=\sum_{i=1}^{n}m_{i}\dot{x}_{i}$, we find that $\dot{\bar{x}}=\bar{v}$. Furthermore, $\dot{\bar{v}}=-\beta$ since the sum of internal forces equals zero by Newton’s third law, meaning that the total momentum changes at a linear rate, while the center of mass will follow a parabolic trajectory. If we instead of the absolute positions $x_{i}$ consider the positions $x_{i}^{\prime}$ relative to the center of mass, i.e., $x_{i}^{\prime}=x_{i}-\bar{x}$, we can consider the system (2.2) without external forces. As we have seen, the external force only affect the center of mass $\bar{x}$, while the internal forces only affect the relative positions $x_{i}^{\prime}$ governed by (2.2) with $x_{i}$ replaced by $x_{i}^{\prime}$ and $\beta=0$. The relative center of mass $\sum_{i=1}^{n}m_{i}x_{i}^{\prime}=0$ and momentum $\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{\prime}=0$ are conserved, and the kinetic energy $E_{\mathrm{k}}=\frac{1}{2}\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{2}$ can be decomposed as
$\frac{1}{2}\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{2}=\frac{1}{2}\dot{\bar{x}}^{2}+\frac{1}{2}\sum_{i=1}^{n}m_{i}(\dot{x}_{i}^{\prime})^{2}.$
@@ 3.1. Elements of optimal transport
(3.2) $d_{\mathscr{W}_{p}}(\rho_{1},\rho_{2})^{2}\coloneqq\min\left\{\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\lvert x-y\rvert^{p}\mathop{}\!\mathrm{d}\varrho(x,y)\colon\varrho\in\mathscr{P}(\mathbb{R}^{n}\times\mathbb{R}^{n}),\>\varpi^{i}_{\#}\varrho=\rho_{i}\right\},$
@@
For any $\rho\in\mathscr{P}(\mathbb{R})$ we define its right-continuous cumulative distribution function
@@ 3.1.1. Comparison to the particle case
The idea is now, analogous to the discrete particle case, to associate (1.1) to a differential inclusion for the pair $(X,V)$. Here the cone $\mathbb{K}^{n}$ is replaced with the set of optimal transport maps $\mathscr{K}$ defined in (3.4), which can be shown to be a closed, convex cone in $L^{2}(\Omega)$. To handle mass concentration, corresponding to particle collisions in the discrete case, we will recall some results on convex analysis.
@@ 3.2. Elements of convex analysis
$\mathscr{N}_{X}\coloneqq\left\{\Xi\in C([0,1])\colon\Xi\geq 0\enspace\mathrm{in}\enspace[0,1]\enspace\mathrm{and}\enspace\Xi=0\enspace\mathrm{in}\enspace\Omega\setminus\Omega_{X}\right\}.$
@@
(3.11) $\operatorname{P}_{\mathscr{H}_{X}}U=\begin{cases}U,&\text{for a.e.\ }m\in\Omega\setminus\Omega_{X}\\ \frac{1}{m_{r}-m_{l}}\int_{m_{l}}^{m_{r}}U(\omega)\mathop{}\!\mathrm{d}\omega,&\text{for a.e.\ }m\in(m_{l},m_{r}),\text{ a maximal interval of }\Omega_{X}.\end{cases}$
@@ 3.3.1. The Lagrangian force representation
(3.19) $F[X]-\operatorname{P}_{\mathscr{H}_{X}}F[X]\in\partial I_{\mathscr{K}}(X)\enspace\text{for all }X\in\mathscr{K}.$
@@ 3.3.2. Definition, existence and uniqueness
$\partial I_{\mathscr{K}}(X(t_{1}))\subset\partial I_{\mathscr{K}}(X(t_{2})),\quad\mathscr{H}_{X(t_{2})}\subset\mathscr{H}_{X(t_{1})}\enspace\text{for any }t_{1}\leq t_{2}.$
@@
$V(t)\in\mathscr{H}_{X(t)}\enspace\text{for all times }t\geq 0$
@@ 3.4. Lagrangian solutions for the Euler system
where, recalling (3.3), $\varrho$ is the unique optimal transport map between the measures $\rho_{1}$ and $\rho_{2}$. The monotone rearrangements of the measures are $X_{\rho_{1}}$ and $X_{\rho_{2}}$, and defining $V_{\rho_{i}}=v_{i}\circ X_{i}$ for $i=1,2$ we observe by (3.3) that we in fact have the relation
@@
A map $f\colon\mathscr{P}_{2}(\mathbb{R})\to\mathscr{M}(\mathbb{R})$ as in 1.2 is bounded if there is a constant $C\geq 0$ such that
@@
(3.27) $d_{\mathscr{T}_{2}}((\rho_{1},f_{\rho_{1}}),(\rho_{2},f_{\rho_{2}}))\leq d_{\mathscr{W}_{2}}(\rho_{1},\rho_{2})\text{ for all }\rho_{1},\rho_{2}\in\mathscr{P}_{2}(\mathbb{R}).$
@@ 3.4.2. Euler–Poisson system with quadratic confinement
A more general force distribution compared to the above is given in [7, Example 6.10]. Given $\sigma\in L^{\infty}(\mathbb{R})$ they define $f[\rho]=-\rho\partial_{x}q_{\sigma}$ for all $\rho\in\mathscr{P}_{2}(\mathbb{R})$, where $q_{\sigma}$ satisfies $-\partial_{xx}q_{\sigma}=\lambda(\rho-\sigma)$. Of course, this only defines $\partial_{x}q_{\sigma}$ up to a constant, and making the same choice for the $\rho$-part as the previous example and defining $S_{\sigma}(x)=\int_{0}^{x}\sigma(y)\mathop{}\!\mathrm{d}y$ one arrives at the following Lagrangian representation
@@ 4.1. Derivation of the scalar conservation law
where we write $M(0,x)=M^{0}(x)$ and $Q(0,x)=Q^{0}(x)$. We know from our assumptions on $\rho_{0}$ and $v_{0}$ that these functions are initially well-defined functions of bounded variation. Then, integrating (1.1a) and (1.1b) together with (1.2), $f[\rho]=f_{\rho}\rho$, yields two advection-type equations
@@
Note that the flux function is exactly the primitive of the prescribed velocity $U$ in (3.20) from Section 3. Differentiating $\mathcal{U}(t,M)$ with respect to $x$ we get a Radon measure which is absolutely continuous with respect to $\rho(t,x)=\partial_{x}M(t,x)$, and the Radon–Nikodym derivative of $\partial_{x}Q$ with respect to $\partial_{x}M$ corresponds to the velocity of the characteristics for the conservation law (4.4), which shows a formal connection between the two solution concepts. If $\rho$ contains atoms, then $M$ contains jumps, and for any sufficiently regular flux function we can apply a $BV$-chain rule as developed in [48]. Note that in general, compositions of $BV$ functions are not necessarily of bounded variation, cf. [29], and in order to have a chain rule for general $BV$ functions we would need the outer function to be at least Lipschitz. However, in our one-dimensional setting where $M$ is the distribution function of a probability measure, we can require less regularity. The following result, found in [36, Lemma 2.1], will be of use.
@@
Consider $p\in[1,\infty]$ and a function $f\in L^{p}(0,1)$, for which we define $F(m)\coloneqq\int_{0}^{m}f(\omega)\mathop{}\!\mathrm{d}\omega$. If $M$ is the right-continuous cumulative distribution function of some Borel probability measure $\mu$ on $\mathbb{R}$, then $F\circ M\in BV(\mathbb{R})$ with distributional derivative $g\mu$, where
(4.6) $g(x)\coloneqq\begin{cases}f\circ M(x),&\mu({x})=0,\\ \frac{F\circ M(x)-F\circ M(x-)}{\mu({x})},&\mu({x})\neq 0\end{cases}$
@@
$\operatorname*{T.\!V.}(f)\coloneqq\sup_{x_{i}}\sum_{i}\lvert f(x_{i})-f(x_{i-1})\rvert<\infty.$
@@
Here $\gamma\in\mathbb{R}$ should satisfy $\gamma>0$ to be considered a damping term. As before, the idea is to integrate (4.10) and consider the cumulative distributions (4.1) of $\rho$ and $\rho v$. Analogous to (4.8) we then arrive at the transport-like equations

### v3

@@ 1. Introduction
with initial density $\rho(0,x)=\rho_{0}(x)$ and velocity $v(0,x)=v_{0}(x)$, where $f[\rho]$ is a force field generated by the fluid itself. Furthermore, we assume initial data $\rho_{0}\in\mathscr{P}_{2}(\mathbb{R})$, the set of probability measures $\mathscr{P}(\mathbb{R})$ with bounded second moments, and $v_{0}\in L^{2}(\mathbb{R},\rho_{0})$, such that the initial kinetic energy is finite. That is, we have
(1.2) $\rho_{0}\in\mathscr{P}_{2}(\mathbb{R})\iff\int_{\mathbb{R}}\lvert x\rvert^{2}\mathop{}\!\mathrm{d}\rho_{0}<\infty\quad\text{and}\quad v_{0}\in L^{2}(\mathbb{R},\rho_{0})\iff\int_{\mathbb{R}}v_{0}^{2}\mathop{}\!\mathrm{d}\rho_{0}<\infty.$
Moreover, we assume the map $f[\rho]\colon\mathscr{P}_{2}(\mathbb{R})\to\mathscr{M}(\mathbb{R})$, where $\mathscr{M}(\mathbb{R})$ are the signed Borel measures with finite total variation, is absolutely continuous with respect to $\rho$, that is,
@@
We are particularly interested in the two, seemingly distinct, notions of solution introduced in [7] and [36, 37]. Each of which yields a unique solution to their respective systems under appropriate assumptions on the forcing terms, even faced with mass concentration. Our aim is to show that these notions in fact are equivalent. More precisely, we prove the equivalence of Lagrangian solutions [7] and entropy solutions [36, 37] for the Euler system (1.1), which contains the attractive and repulsive Euler–Poisson systems (1.4) as special cases. A very recent paper [43] describes this equivalence between sticky particles flow described in [28] and entropy solutions of (1.1) for $f[\rho]\equiv 0$ .
@@
with flux function $\mathcal{A}(m)=\alpha(m-m^{2})$ and $M$ being the (cumulative) distribution function of $\mu$, $M(t,x)=\mu(t,(-\infty,x])$. Here $\alpha=1$ results in an attractive potential, while $\alpha=-1$ gives a repulsive potential. The dynamics of (1.5) is relatively simple in the sense that an attractive potential promotes accumulation of mass $\mu$, while a repulsive potential spreads the mass. This is particularly clear for a sum of Dirac masses $\mu=\sum_{i=1}^{n}m_{i}\delta_{x_{i}}$, i.e., a set of $n$ “particles”, as an attractive potential will aggregate the particles into one particle of unit mass, while a repulsive potential will instantly diffuse the Dirac measures into a measure which is absolutely continuous with respect to the Lebesgue measure. The aggregation and diffusion turn out to have a one-to-one correspondence with respectively shock and rarefaction waves of the conservation law (1.6), see [4, Section 2]. In fact, there the equivalence of (1.5) and (1.6) is first established at the particle level before passing to a limit to cover the general case.
@@
In the above, $\gamma\geq 0$ is a linear damping parameter, while $\tilde{\phi}=\frac{1}{2}x^{2}-\lvert x\rvert$ represents a potential with Coulomb repulsion and quadratic confinement. This system has been studied in detail in [14] for classical solutions up until blow-up, or mass concentration, for which precise critical thresholds are given. We want to use the aforementioned solution concepts to study (1.7) beyond the time of blow-up and its asymptotic behavior.
@@ 1.1. Main results
(1.9) $\frac{v(t,x_{2})-v(t,x_{1})}{x_{2}-x_{1}}\leq\frac{1}{t}\kern 5.0pt\text{for $\rho$-a.e.}\ x_{1}\leq x_{2}.$
@@
Concept $\backslash$ Framework Lagrangian solution Entropy solution
@@ 2.1. First considerations and simplifications
The first term on the right-hand side of (2.2), scaled by $\alpha$, is the force per unit mass coming from the Poisson interaction forces between particles, and is therefore an internal force for the system. On the other hand, the second term scaled by $\beta$ is the force per unit mass coming from an external force, which could be an external gravitational or electrical field. Introducing the kinetic energy $E_{\mathrm{k}}(\{x_{i}\})=\frac{1}{2}\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{2}$, then the Euler–Lagrange equations for the Lagrangian $L=E_{\mathrm{k}}-E_{\mathrm{p}}$ yields (2.2). Moreover, the energy $E_{\mathrm{k}}+E_{\mathrm{p}}$, which corresponds to the Hamiltonian after introducing the momentum $p_{i}=m_{i}\dot{x}_{i}$, is conserved. The total momentum $\sum_{i=1}^{n}m_{i}\dot{x}_{i}$ however, is not conserved because of the external force. If we introduce the center of mass $\bar{x}\coloneqq\sum_{i=1}^{n}m_{i}x_{i}$ and average velocity, which here coincides with the total momentum, $\bar{v}=\sum_{i=1}^{n}m_{i}\dot{x}_{i}$, we find that $\dot{\bar{x}}=\bar{v}$. Furthermore, $\dot{\bar{v}}=-\beta$ since the sum of internal forces equals zero by Newton’s third law, meaning that the total momentum changes at a linear rate, while the center of mass will follow a parabolic trajectory. If we instead of the absolute positions $x_{i}$ consider the positions $x_{i}^{\prime}$ relative to the center of mass, i.e., $x_{i}^{\prime}=x_{i}-\bar{x}$, we can consider the system (2.2) without external forces. As we have seen, the external force only affect the center of mass $\bar{x}$, while the internal forces only affect the relative positions $x_{i}^{\prime}$ governed by (2.2) with $x_{i}$ replaced by $x_{i}^{\prime}$ and $\beta=0$. The relative center of mass $\sum_{i=1}^{n}m_{i}x_{i}^{\prime}=0$ and momentum $\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{\prime}=0$ are conserved, and the kinetic energy $E_{k}$ can be decomposed as
$E_{\mathrm{k}}(\{\dot{x}_{i}\})=\frac{1}{2}\sum_{i=1}^{n}m_{i}\dot{x}_{i}^{2}=\frac{1}{2}\dot{\bar{x}}^{2}+\frac{1}{2}\sum_{i=1}^{n}m_{i}(\dot{x}_{i}^{\prime})^{2}.$
@@ 3.1. Elements of optimal transport
(3.2) $d_{\mathscr{W}_{p}}(\rho_{1},\rho_{2})^{p}\coloneqq\min\left\{\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\lvert x-y\rvert^{p}\mathop{}\!\mathrm{d}\varrho(x,y)\colon\varrho\in\mathscr{P}(\mathbb{R}^{n}\times\mathbb{R}^{n}),\>\varpi^{i}_{\#}\varrho=\rho_{i}\right\},$
@@
For any $\rho\in\mathscr{P}(\mathbb{R})$ we define its right-continuous distribution function
@@ 3.1.1. Comparison to the particle case
The idea is now, analogous to the discrete particle case, to associate (1.1) to a differential inclusion for the pair $(X,V)$. Here the cone $\mathbb{K}^{n}$ is replaced with the set of optimal transport maps $\mathscr{K}$ defined in (3.4), which can be shown to be a closed, convex cone in $L^{2}(\Omega)$. To handle mass concentration, corresponding to particle collisions in the discrete case, we will recall some results from convex analysis.
@@ 3.2. Elements of convex analysis
$\mathscr{N}_{X}\coloneqq\left\{\Xi\in C([0,1])\colon\Xi\geq 0\kern 5.0pt\mathrm{in}\kern 5.0pt[0,1]\kern 5.0pt\mathrm{and}\kern 5.0pt\Xi=0\kern 5.0pt\mathrm{in}\kern 5.0pt\Omega\setminus\Omega_{X}\right\}.$
@@
[…]
@@ 3.3.1. The Lagrangian force representation
(3.19) $F[X]-\operatorname{P}_{\mathscr{H}_{X}}F[X]\in\partial I_{\mathscr{K}}(X)\kern 5.0pt\text{for all }X\in\mathscr{K}.$
@@ 3.3.2. Definition, existence and uniqueness
$\partial I_{\mathscr{K}}(X(t_{1}))\subset\partial I_{\mathscr{K}}(X(t_{2})),\quad\mathscr{H}_{X(t_{2})}\subset\mathscr{H}_{X(t_{1})}\kern 5.0pt\text{for any }t_{1}\leq t_{2}.$
@@
$V(t)\in\mathscr{H}_{X(t)}\kern 5.0pt\text{for all times }t\geq 0$
@@ 3.4. Lagrangian solutions for the Euler system
where, recalling (3.3), $\varrho$ is the unique optimal transport map between the measures $\rho_{1}$ and $\rho_{2}$. Here we may recall our assumptions (1.2) on the initial data $(\rho_{0},v_{0})$ and observe that this aligns with the definition of (3.24). The monotone rearrangements of the measures are $X_{\rho_{1}}$ and $X_{\rho_{2}}$, and defining $V_{\rho_{i}}=v_{i}\circ X_{i}$ for $i=1,2$ we observe by (3.3) that we in fact have the relation
@@
A map $f\colon\mathscr{P}_{2}(\mathbb{R})\to\mathscr{M}(\mathbb{R})$ as in 1.3 is bounded if there is a constant $C\geq 0$ such that
@@
(3.27) $d_{\mathscr{T}_{2}}((\rho_{1},f_{\rho_{1}}),(\rho_{2},f_{\rho_{2}}))\leq\omega\left(d_{\mathscr{W}_{2}}(\rho_{1},\rho_{2})\right)\text{ for all }\rho_{1},\rho_{2}\in\mathscr{P}_{2}(\mathbb{R}).$
@@ 3.4.2. Euler–Poisson system with quadratic confinement
A more general force distribution compared to the above is given in [7, Example 6.10]. Given $\sigma\in L^{\infty}(\mathbb{R})$ they define $f[\rho]=-\rho\partial_{x}q_{\sigma}$ for all $\rho\in\mathscr{P}_{2}(\mathbb{R})$, where $q_{\sigma}$ satisfies $\partial_{xx}q_{\sigma}=\lambda(\rho-\sigma)$. Of course, this only defines $\partial_{x}q_{\sigma}$ up to a constant, and making the same choice for the $\rho$-part as the previous example and defining $S_{\sigma}(x)=\int_{0}^{x}\sigma(y)\mathop{}\!\mathrm{d}y$ one arrives at the following Lagrangian representation
@@ 4.1. Derivation of the scalar conservation law
where we have used the dummy variable $y$ to indicate integration over the spatial variable. In order to clarify the notation in the rest of the paper, we emphasize that integration in space with respect to a time-dependent measure $\rho=\rho(t,x)$ is indicated by the differential $\mathop{}\!\mathrm{d}\rho(t,\cdot)$, where we here have omitted the variable of integration. If there is additional integration in time, there will be a separate, additional differential for this, e.g., $\mathop{}\!\mathrm{d}t$. We write $M(0,x)=M^{0}(x)$ and $Q(0,x)=Q^{0}(x)$ for the initial data of (4.1), and we recall from our assumptions (1.2) on $\rho_{0}$ and $v_{0}$ that these distribution functions are initially well-defined functions of bounded variation. Then, integrating (1.1a) and (1.1b) together with (1.3), $f[\rho]=f_{\rho}\rho$, yields two advection-type equations
@@
Note that the flux function $\mathcal{U}$ is exactly the primitive of the prescribed velocity $U$ in (3.20) from Section 3. Differentiating $\mathcal{U}(t,M)$ with respect to $x$ we get a Radon measure which is absolutely continuous with respect to $\rho(t,x)=\partial_{x}M(t,x)$, and the Radon–Nikodym derivative of $\partial_{x}Q$ with respect to $\partial_{x}M$ corresponds to the velocity of the characteristics for the conservation law (4.4), which shows a formal connection between the two solution concepts. If $\rho$ contains atoms, then $M$ contains jumps, and for any sufficiently regular flux function we can apply a $BV$-chain rule as developed in [48]. Note that in general, compositions of $BV$ functions are not necessarily of bounded variation, cf. [29], and in order to have a chain rule for general $BV$ functions we would need the outer function to be at least Lipschitz. However, in our one-dimensional setting where $M$ is the distribution function of a probability measure, we can require less regularity. The following result, found in [36, Lemma 2.1], will be of use.
@@
Consider $p\in[1,\infty]$ and a function $f\in L^{p}(0,1)$, for which we define $F(m)\coloneqq\int_{0}^{m}f(\omega)\mathop{}\!\mathrm{d}\omega$. If $M$ is the right-continuous distribution function of some Borel probability measure $\mu$ on $\mathbb{R}$, then $F\circ M\in BV(\mathbb{R})$ with distributional derivative $g\mu$, where
(4.6) $g(x)\coloneqq\begin{cases}f\circ M(x),&\mu(\{x\})=0,\\ \frac{F\circ M(x)-F\circ M(x-)}{\mu(\{x\})},&\mu(\{x\})\neq 0\end{cases}$
@@
$\operatorname*{T.\!V.}(f)\coloneqq\sup_{x_{i}}\sum_{i}\lvert f(x_{i})-f(x_{i-1})\rvert$$\infty$.
@@
Here $\gamma\in\mathbb{R}$ should satisfy $\gamma>0$ to be considered a damping term. As before, the idea is to integrate (4.10) and consider the distribution functions (4.1) of $\rho$ and $\rho v$. Analogous to (4.8) we then arrive at the transport-like equations